OpenStudy (anonymous):

At a club meeting, there are 10 more club members than nonmembers. The number of club members at the meeting is c. Quantity A: The total number of people at the club meeting. Quantity B: 2c-10

OpenStudy (anonymous):

The two quantities are equal... is this the answer your looking for?

OpenStudy (anonymous):

i have the answer already i need to know how to do it

OpenStudy (anonymous):

I don't understand how theyre equal because when you use an imaginary number Quantity B is larger

OpenStudy (anonymous):

(10)+10=20

OpenStudy (anonymous):

2(20)-10= 30

OpenStudy (anonymous):

C + C-10 = 2C-10

OpenStudy (anonymous):

@ganeshie8